OpenStudy (gigi1103):

What is the equation in point-slope form of a line that passes through the points (5, −3) and (−2, 9) ? ive gotten to 12/-7

OpenStudy (gigi1103):

@princeharryyy

jabez177 (jabez177):

y – y1 = m(x – x1)

OpenStudy (gigi1103):

y-9=idk*(x- -2)

jabez177 (jabez177):

Ugh, sorry! I lost track... I forgot how to do this again...

OpenStudy (gigi1103):

@EclipsedStar

OpenStudy (gigi1103):

@Vocaloid

OpenStudy (harman.singh):

I gave this question a go and got the same gradient you mentioned in your question ie -12/7 y-9=-12/7(x+2)

OpenStudy (gigi1103):

thank you!

OpenStudy (princeharryyy):

SORRY! i GOT OFFLINE. @gigi1103

OpenStudy (gigi1103):

its ok!

OpenStudy (princeharryyy):

TNK U

OpenStudy (princeharryyy):

this idk, y-9=idk*(x- -2) is the slope(m) that u calculate by youself. @gigi1103